# Securing the Skies: A Comprehensive Survey on Anti-UAV Methods, Benchmarking, and Future Directions

### v2

@@ 3 UAV Classification Approaches
Table 2 synthesizes representative UAV classification methods covering radio frequency (RF), audio, radar, vision, and LiDAR modalities. These methods range from simple drone-presence detection [87] to more nuanced classifications of UAV types, flight modes, and orientations [109]. Although comparison is difficult due to inconsistent experimental protocols, modality-specific considerations—such as spectral noise or lighting variability—motivate distinct architectural and data-processing strategies. Below, we highlight key insights and emerging trends.
@@
(1) RF-Based Classification. RF-focused approaches exploit the electromagnetic signatures emitted by UAVs to detect, identify, and even recognize specific flight modes. Al-Sa’d et al. [3] introduced the DroneRF dataset, demonstrating high accuracy (99.7%) on a 2-class task but lower performance (46.8%) on a more granular 10-class task. Subsequent work by Al-Emadi et al. [1] confirmed CNNs’ viability for DroneRF, achieving 99.8% (2-class) and 59.2% (10-class). Recent studies have explored advanced feature engineering or ensemble learning: for instance, Nemer et al. [86] outperformed CNN-based baselines via hierarchical ensembles, while Li et al. [72] incorporated bispectrum representations with contrastive learning. Inani et al. [56] achieved near-perfect 10-class accuracy (99.51%) by fusing PSD features, a 1D-CNN, and XGBoost. Despite these advancements, real-world deployments face challenges including spectral congestion, adversarial jamming, and domain shifts [9, 37, 132].
@@
(2) Audio-Based Classification. Audio pipelines distinguish UAVs based on rotor noise and other acoustic signatures. Al-Emadi et al. [2] introduced the DroneAudio dataset, reporting 96.38% (2-class) and 92.94% (3-class) accuracy using CNNs. Katta et al. [61] subsequently demonstrated the benefits of LSTM, CRNN, and Transformers, with LSTM attaining 98.93% (2-class). Meanwhile, Xiao et al. [148] employed a temporal CNN+RNN to handle MMAUD’s audio modality, reaching 98.0% accuracy alongside strong pose estimation results. Noise from wind, background chatter, or multiple UAVs remains an obstacle in practical scenarios [14, 57], motivating research into denoising filters and multi-microphone arrays [119].
@@
(3) Radar-Based Classification. Radar offers micro-Doppler signatures that persist under reduced visibility or nighttime conditions. Early studies (e.g., Mendis et al. [81]) validated radar-based UAV classification using deep belief networks, though limited dataset availability inhibited broad adoption. CNN- and LSTM-based pipelines now dominate, focusing on micro-Doppler representations to improve robustness in cluttered environments [166, 132]. Nonetheless, radar hardware cost and calibration constraints remain practical challenges [132].
@@
(4) Vision-Based Classification. Computer vision pipelines target UAV recognition via optical images, often focusing on manufacturer types (e.g., DJI Phantom, Mavic). Wisniewski et al. [139] achieved 92.4% accuracy on three drone types (Phantom, Mavic, Inspire) and illustrated synthetic imagery’s viability [140]. Other studies address issues of small object size and motion blur, incorporating advanced detectors or domain adaptation [130, 109]. While performance in controlled scenarios is high, vision-based methods degrade under extreme illumination, adverse weather, or dense clutter [18, 33].
@@
(5) Multi-Modal Recognition. Integrating multiple sensors (e.g., RF+Audio [43], Radar+LiDAR+Vision [156], or Audio+Vision [147]) can mitigate single-sensor blind spots. Deng et al. [32] fused fisheye RGB, radar, and LiDAR data from MMAUD, reaching 81.36% accuracy on a 4-class task, while Xiao et al. [147] combined audio and RGB streams to achieve 99.3% accuracy and sub-meter pose errors. Although multi-modal setups improve robustness, they also require careful calibration, synchronization, and greater computational resources [156, 132].
@@
Key Insights & Challenges. Across these approaches [33, 166, 132, 130, 18], each modality faces inherent constraints: (1) RF-based solutions excel at flight-mode or type identification [3, 9], but radio congestion and potential adversarial attacks are critical hurdles [37, 132]. (2) Audio-based methods achieve high accuracy in controlled settings yet suffer from significant performance drops in high-noise or windy environments [14, 57, 2]. (3) Radar-based pipelines bypass visual occlusions but incur higher costs and calibration demands [81, 132]. (4) Vision-based approaches generally provide fine-grained recognition but degrade with lighting or weather changes [139, 130]. (5) Multi-modal fusion can mitigate single-sensor weaknesses and reduce false positives [32, 43] yet raises complexity in real-time deployment [156]. Additionally, adversarial robustness remains an open problem, particularly where sensors or neural networks may be fooled by spoofing or deceptive signals [80, 61]. A continuing research direction involves efficient, low-latency architectures [148, 132] that can handle on-board or edge-based processing for rapid UAV detection and classification.
@@
Jiang et al. (2023) [58] | RGB, IR | GlobalTrack + DFSC | Anti-UAV [58] | Released 318-sequence RGB-T UAV tracking benchmark. Achieved the best state accuracy score of 74.33% on the validation set and 65.41% on the test set |
Wu et al. (2024) [144] | RGB | DCF + dual-fovea + contrast + opponent-color (EECF) | Anti-UAV [58], USC-Drone [24] | Real-time EECF tracker (86 FPS) via dual-fovea design. Achieved the state accuracy score of 66.42% on Anti-UAV, 70.45% on USC-Drone. |
Wang et al. (2024) [134] | RGB | YOLOX + Swin-T + -IoU + BYTE | UAVSwarm [133] | Detection-driven approach with -IoU + ByteTrack. Achieved 78.2% tracking accuracy on UAVSwarm test split. |
Deng et al.(2024) [32] | RGB, RA, L | YOLOv9 + EfficientNet-B7 + LSTM | MMAUD [156] | Intergrated YOLOv9 + EfficientNet-B7 + LSTM/Kalman for classification and tracking. Achieved 2.21 Pose MSE Loss on 3D tracking of MMAUD and ranked 1st in UG2+ Challenge. |
@@
Huang et al. (2021) [49] | IR | SiamSTA | 1st / 2nd Anti-UAV Challenges [58] [162] | Achieved 74.46 and 67.30 scores on 1st / 2nd Anti-UAV test-dev sets, ranking 1st. Extended SiamSTA with spatio-temporal attention + motion estimation. |
Li et al. (2022) [71] | IR | SuperDiMP (ResNet50) | 2nd Anti-UAV Challenge [162] | Used SuperDiMP with dual-branch IR fusion. Achieved +5.84% SR and +4.52% PR over baseline. Focused on contrast inversion robustness. |
Chen et al. (2022) [22] | IR | SiamSTA + CDCF | 1st Anti-UAV Challenge [58] | SiamSTA with CDCF and 3-stage redetection. Achieved top precision in 2nd Anti-UAV Challenge. Achieved 68.26 and 76.11 scores on Anti-UAV test and validation sets. Robust on fast/small UAVs. |
Zhang et al. (2023) [161] | RGB, IR | SiamFusion | Anti-UAV [58] | Achieved overall 63.37 score on Anti-UAV with 14 FPS. Developed a dual-fusion tracker with modality + decision-level fusion. Used Swin backbone + Local-Global Converter. |
Xie et al. (2023) [149] | IR | FPN-based Siamese + STG-RPN + ID-RCNN (STFTrack) | Anti-UAV [58] | STFTrack using STG-RPN and ID-RCNN. Achieved 91.2% precision, 66.6% success @ 12.4 FPS. Addressed distractors and thermal crossover. |
Zhang et al. (2024) [158] | IR | OSTrack + MCJT (RMCM + FSRM + ATUS ) | 1st / 2nd / 3rd Anti-UAV Challenges [58] [162] [164] | Achieved 69.65% Acc, 80.45% Acc, 65.88% Acc and 50.28% Acc on the 1st Anti-UAV test set, 1st Anti-UAV test-dev, 2nd Anti-UAV test-dev, and 3rd Anti-UAV validation set. Proposed MCJT (RMCM + FSRM + ATUS) with adaptive template and multi-consistency. Ran at 44 FPS. |
Huang et al. (2024) [50] | IR | SiamSRT + C-C RPN + S-L RCNN + TMB + SCFD | Anti-UAV [58] | Achieved 71.64 and 80.04 scores on Anti-UAV testing and validation sets. Global Siamese tracker using Swin-T + consistency modules (TMB + SCFD). Avoided template degradation; outperformed multiple SOTA trackers. |
Huang et al. (2024) [51] | IR | SiamDT (Swin Transformer + Dual-Semantic RPN + Versatile R-CNN) | Anti-UAV410 [51] | Introduced Anti-UAV410 dataset. Proposed SiamDT with dual semantic reasoning + background suppression. Achieved 68.19 and 71.65 scores on its testing and validation sets. |
@@
Cheng et al. (2022) [25] | IR | Multi-Hybrid Attention + SiamRPN++ + YOLOv5 (SiamAD) | Anti-UAV [58] | SiamAD with CSAM/CAM + hierarchical discriminator + YOLOv5. 88.4% DP @ 37.1 FPS. Improved long-term IR tracking. |
Yu et al. (2023) [155] | IR | Transformer + YOLOv5 + LoFTR + GMM + OSTrack | 1st / 2nd Anti-UAV Challenges [58] [162] | Achieved 2nd place in Anti-UAV Challenge (3rd) with 68.8 score. And 75.13% and 80.38% accuracy on the 1st and 2nd test dev, individually. Integrated OSTrack, YOLOv5, LoFTR, GMM for multi-stage IR tracking. |
@@ 4 UAV Detection Approaches
(1) Radar-Based Detection. Radar methods detect UAVs by exploiting micro-Doppler signatures, providing resilience under low-visibility or harsh weather conditions [165, 132]. For instance, Zhao et al. [165] achieved decimeter-level localization at 10 Hz by combining CNNs with LSTMs, outperforming conventional radar baselines. However, radar sensors can suffer from interference and clutter, particularly in urban areas or heavily congested electromagnetic environments [143].
@@
(2) Vision-Based Detection. Vision-based techniques remain widely adopted, benefiting from readily available cameras and sophisticated CNN-based detectors [170, 124]. Early work centered on region-based methods [99] and YOLO architectures [17, 135], gradually evolving into more lightweight or attention-oriented variants [26, 53, 46]. While mosaic augmentation and specialized modules can improve small-object detection [31, 26], vision-based approaches remain vulnerable to occlusions, clutter, and extreme distances. Researchers have thus explored thermal and infrared (IR) imaging [118, 85] or speed-accuracy trade-offs for real-time embedded scenarios [16, 38].
@@
(3) Multi-Modal Fusion. Single-modality approaches often fail under suboptimal conditions—low-visibility (vision), interference (radar), or noise (audio) [2, 129]. Consequently, multi-modal fusion has gained traction as a means of leveraging complementary sensor data [120, 12]. Svanström et al. [120] showed that augmenting YOLO-based RGB or thermal imaging with audio features increased F1 scores and reduced false alarms. Similarly, Wu et al. [143] fused radar and RGB streams via attention-based spatiotemporal networks to improve occlusion handling. Practical adoption, however, demands meticulous sensor synchronization, balanced hardware costs, robust computational efficiency, and real-time data association [98].
@@
Key Insights & Challenges. From prior work [132, 170, 124], several recurring themes emerge. 1) Small UAV Detection: Limited pixel coverage and subtle micro-Doppler signatures complicate detection [47, 20, 77], prompting multi-scale feature extraction, deformable convolutions, and transformer backbones [124, 26, 53]. Ensuring real-time operation at high accuracy remains challenging [132, 135]. 2) Environmental Factors: Ambient noise (audio), lighting (vision), and radio interference (radar) degrade performance [2, 129, 113, 165], highlighting the need for advanced clutter suppression and robust sensor hardware. 3) Generalization vs. Specialization: YOLO-based models offer generalizable frameworks [31, 26, 46], but specialized domains (e.g., stealth UAVs) may require niche architectures or domain-adaptive strategies [143, 16]. 4) Multi-Modal Fusion: Complementary sensors effectively reduce false positives and enhance accuracy [120, 143], although issues like sensor alignment, real-time latency, and hardware cost must be resolved [12, 98]. 5) Benchmarking & Deployment: Heterogeneous data sources, labeling formats, and test conditions [35, 93, 170] complicate standardized evaluations. Though laboratory demonstrations show promise, real-time performance in crowded outdoor settings remains largely underexplored [39, 42]. Overcoming these challenges will be crucial for field-ready anti-UAV systems.
forked edges,
for tree=
grow=east,
reversed=true,
anchor=base west,
parent anchor=east,
child anchor=west,
base=center,
font=,
rectangle,
draw=gray,
rounded corners,
align=left,
text centered,
minimum width=4em,
edge+=darkgray, line width=1pt,
s sep=3pt,
inner xsep=2pt,
inner ysep=3pt,
line width=0.8pt,
ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center,
where level=1text width=18em,font=,,
where level=2text width=18em,font=,,
where level=3text width=18em,font=,,
where level=4text width=35em,font=,,
where level=5text width=12em,font=,,
Emerging Methods and Technologies, ver, line width=0.7mm,fill=gray!20, text width=25em
Diffusion Models for Data
Synthesis and Domain Adaptation
, fill=c1!60, draw=c1, line width=0mm
Foundational Theories, fill=c1!60, draw=c1, line width=0mm, edge=c1
Diffusion Model,fill=c1!60, draw=c1, line width=0mm, edge=c1
Sohl-Dickstein et al. (2015) [115]; Ho et al. (2020) [48];
Song and Ermon (2021) [116], leaf, draw=c1, line width=0.7mm, text width=28em, edge=c1
Data Synthesis,fill=c1!60, draw=c1, line width=0mm, edge=c1
Lu et al. (2023) [79], leaf, draw=c1, line width=0.7mm, text width=28em, edge=c1
Domain Adaptation,fill=c1!60, draw=c1, line width=0mm,edge=c1
Pan and Yang (2010) [91]; Torralba and Efros (2011) [126];
Tsai et al. (2018) [127], leaf, draw=c1, line width=0.7mm, text width=28em, edge=c1
Text-Conditioned Synthesis, fill=c1!60, draw=c1, line width=0mm,edge=c1
Ramesh et al. (2022) [97]; Saharia et al. (2022) [105], leaf, draw=c1, line width=0.7mm, text width=47.5em, edge=c1
High-Resolution Synthesis, fill=c1!60, draw=c1, line width=0mm,edge=c1
Rombach et al. (2022) [101], leaf, draw=c1, line width=0.7mm, text width=47.5em, edge=c1
Efficient Synthesis Generation, fill=c1!60, draw=c1, line width=0mm,edge=c1
Salimans and Ho (2022) [106]; Qiao et al. (2025) [95], leaf, draw=c1, line width=0.7mm, text width=47.5em, edge=c1
Unsupervised Domain Adaptation, fill=c1!60, draw=c1, line width=0mm, edge=c1
Peng et al. (2024) [94]; Benigmim et al. (2023) [13]; Zhang et al. (2025) [160], leaf, draw=c1, line width=0.7mm, text width=47.5em, edge=c1
Vision-Language Models
and Semantic Context
, fill=c2!60, draw=c2, line width=0mm
Pre-training, fill=c2!60, draw=c2, line width=0mm, edge=c2
Su et al. (2019) [117]; Radford et al. (2021) [96]; Zhong et al. (2022) [167]; Li et al. (2022) [70];
Alayrac et al. (2022) [5]; Li et al. (2023) [69]; Huang et al. (2023) [54], leaf, draw=c2, line width=0.7mm, text width=47.5em, edge=c2
Open-Vocabulary Detection, fill=c2!60, draw=c2, line width=0mm, edge=c2
Kamath et al. (2021) [60]; Zareian et al. (2021) [157]; Gu et al. (2022) [45]; Zhou et al. (2022) [168], leaf, draw=c2, line width=0.7mm, text width=47.5em, edge=c2
UAV Pioneers, fill=c2!60, draw=c2, line width=0mm, edge=c2
Detection, fill=c2!60, draw=c2, line width=0mm, edge=c2
Liu et al. (2024) [76]; Li et al. (2024) [73], leaf, draw=c2, line width=0.7mm, text width=28em, edge=c2
Semantic Segmentation, fill=c2!60, draw=c2, line width=0mm, edge=c2
Huang et al. (2025) [52], leaf, draw=c2, line width=0.7mm, text width=28em, edge=c2
Tracking, fill=c2!60, draw=c2, line width=0mm, edge=c2
Feng et al. (2024) [41], leaf, draw=c2, line width=0.7mm, text width=28em, edge=c2
Self-Supervised and Unsupervised
Learning
, fill=c3!60, draw=c3, line width=0mm
Domain Adaptation, fill=c3!60, draw=c3, line width=0mm, edge=c3
Kim et al. (2019) [64]; Kim et al. (2019) [65]; Ye et al. (2022) [152], leaf, draw=c3, line width=0.7mm, text width=47.5em, edge=c3
UAV Pioneers, fill=c3!60, draw=c3, line width=0mm, edge=c3
Security, fill=c3!60, draw=c3, line width=0mm, edge=c3
Alanazi (2024) [4], leaf, draw=c3, line width=0.7mm, text width=28em, edge=c3
Swarm, fill=c3!60, draw=c3, line width=0mm, edge=c3
Mou et al. (2024) [84], leaf, draw=c3, line width=0.7mm, text width=28em, edge=c3
Detection, fill=c3!60, draw=c3, line width=0mm, edge=c3
Kim et al. (2023) [63], leaf, draw=c3, line width=0.7mm, text width=28em, edge=c3
Tracking, fill=c3!60, draw=c3, line width=0mm, edge=c3
Lai et al. (2020) [66]; Sio et al. (2020) [114];
Wang et al. (2021) [137]; Cao et al. (2021) [19], leaf, draw=c3, line width=0.7mm, text width=28em, edge=c3
Recognition, fill=c3!60, draw=c3, line width=0mm, edge=c3
Lu et al. (2024) [78];, leaf, draw=c3, line width=0.7mm, text width=28em, edge=c3
Multi-Modal Fusion and
Sensor Integration
, fill=c4!60, draw=c4, line width=0mm
Multi-Sensor Foundations, fill=c4!60, draw=c4, line width=0mm, edge=c4
Samaras et al. (2019) [107]; Svanström et al. (2021) [120]; Dudczyk et al. (2022) [36]; Zhu et al. (2023) [171], leaf, draw=c4, line width=0.7mm, text width=47.5em, edge=c4
Model-driven Fusion, fill=c4!60, draw=c4, line width=0mm, edge=c4
Park et al. (2015) [92]; Liu et al. (2017) [75]; Al-Emadi et al. (2019) [2]; Kumar et al. (2024) [27], leaf, draw=c4, line width=0.7mm, text width=47.5em, edge=c4
Deep Embedded Fusion, fill=c4!60, draw=c4, line width=0mm, edge=c4
Alla et al. (2024) [7]; Yang et al. (2024) [151]; Wu et al. (2024) [143], leaf, draw=c4, line width=0.7mm, text width=47.5em, edge=c4
Reinforcement Learning and
Autonomous Control
, fill=c5!60, draw=c5, line width=0mm
Foundational Theories, fill=c5!60, draw=c5, line width=0mm, edge=c5
Mnih et al. (2015) [83]; Silver et al. (2016) [112]; Lillicrap et al. (2015) [74]; Sadeghi and Levine (2016) [104], leaf, draw=c5, line width=0.7mm, text width=47.5em, edge=c5
UAV Pioneers, fill=c5!60, draw=c5, line width=0mm, edge=c5
Interception and Countermeasures, fill=c5!60, draw=c5, line width=0mm, edge=c5
Çetin et al. (2020) [172]; Bertoin et al. (2022) [15];
Çetin et al. (2022) [173], leaf, draw=c5, line width=0.7mm, text width=28em, edge=c5
Detection, fill=c5!60, draw=c5, line width=0mm, edge=c5
AlKhonaini et al. (2024) [6], leaf, draw=c5, line width=0.7mm, text width=28em, edge=c5
Tracking, fill=c5!60, draw=c5, line width=0mm, edge=c5
Wisniewski et al. (2024) [142], leaf, draw=c5, line width=0.7mm, text width=28em, edge=c5
Drone Racing, fill=c5!60, draw=c5, line width=0mm, edge=c5
Kaufmann et al. (2023) [62], leaf, draw=c5, line width=0.7mm, text width=28em, edge=c5

### v3

@@ 3 UAV Classification Approaches
Table 2 synthesizes representative UAV classification methods covering radio frequency (RF), audio, radar, vision, and LiDAR modalities. These methods range from simple drone-presence detection [Nguyen2017] to more nuanced classifications of UAV types, flight modes, and orientations [Scholes2022]. Although comparison is difficult due to inconsistent experimental protocols, modality-specific considerations—such as spectral noise or lighting variability—motivate distinct architectural and data-processing strategies. Below, we highlight key insights and emerging trends.
@@
(1) RF-Based Classification. RF-focused approaches exploit the electromagnetic signatures emitted by UAVs to detect, identify, and even recognize specific flight modes. Al-Sa’d et al. [AlSaad2019] introduced the DroneRF dataset, demonstrating high accuracy (99.7%) on a 2-class task but lower performance (46.8%) on a more granular 10-class task. Subsequent work by Al-Emadi et al. [AlEmadi2020cnn] confirmed CNNs’ viability for DroneRF, achieving 99.8% (2-class) and 59.2% (10-class). Recent studies have explored advanced feature engineering or ensemble learning: for instance, Nemer et al. [nemer2021] outperformed CNN-based baselines via hierarchical ensembles, while Li et al. [li2021bissiam] incorporated bispectrum representations with contrastive learning. Inani et al. [Inani2023] achieved near-perfect 10-class accuracy (99.51%) by fusing PSD features, a 1D-CNN, and XGBoost. Despite these advancements, real-world deployments face challenges including spectral congestion, adversarial jamming, and domain shifts [Allahham2020, Elleuch2024jamming, wang2024survey].
@@
(2) Audio-Based Classification. Audio pipelines distinguish UAVs based on rotor noise and other acoustic signatures. Al-Emadi et al. [al2019droneaudio] introduced the DroneAudio dataset, reporting 96.38% (2-class) and 92.94% (3-class) accuracy using CNNs. Katta et al. [Katta2022] subsequently demonstrated the benefits of LSTM, CRNN, and Transformers, with LSTM attaining 98.93% (2-class). Meanwhile, Xiao et al. [xiao2025tame] employed a temporal CNN+RNN to handle MMAUD’s audio modality, reaching 98.0% accuracy alongside strong pose estimation results. Noise from wind, background chatter, or multiple UAVs remains an obstacle in practical scenarios [Bernardini2017, Jeon2017], motivating research into denoising filters and multi-microphone arrays [Sun2023DeepLD].
@@
(3) Radar-Based Classification. Radar offers micro-Doppler signatures that persist under reduced visibility or nighttime conditions. Early studies (e.g., Mendis et al. [mendis2016deep]) validated radar-based UAV classification using deep belief networks, though limited dataset availability inhibited broad adoption. CNN- and LSTM-based pipelines now dominate, focusing on micro-Doppler representations to improve robustness in cluttered environments [zheng2021air, wang2024survey]. Nonetheless, radar hardware cost and calibration constraints remain practical challenges [wang2024survey].
@@
(4) Vision-Based Classification. Computer vision pipelines target UAV recognition via optical images, often focusing on manufacturer types (e.g., DJI Phantom, Mavic). Wisniewski et al. [Wisniewski2022] achieved 92.4% accuracy on three drone types (Phantom, Mavic, Inspire) and illustrated synthetic imagery’s viability [Wisniewski2022dataset]. Other studies address issues of small object size and motion blur, incorporating advanced detectors or domain adaptation [Unlu2019, Scholes2022]. While performance in controlled scenarios is high, vision-based methods degrade under extreme illumination, adverse weather, or dense clutter [cabrera2019detection, dewangan2023application].
@@
(5) Multi-Modal Recognition. Integrating multiple sensors (e.g., RF+Audio [Frid2024], Radar+LiDAR+Vision [yuan2024mmaud], or Audio+Vision [xiao2024av]) can mitigate single-sensor blind spots. Deng et al. [deng2024multi] fused fisheye RGB, radar, and LiDAR data from MMAUD, reaching 81.36% accuracy on a 4-class task, while Xiao et al. [xiao2024av] combined audio and RGB streams to achieve 99.3% accuracy and sub-meter pose errors. Although multi-modal setups improve robustness, they also require careful calibration, synchronization, and greater computational resources [yuan2024mmaud, wang2024survey].
@@
Key Insights & Challenges. Across these approaches [dewangan2023application, zheng2021air, wang2024survey, Unlu2019, cabrera2019detection], each modality faces inherent constraints: (1) RF-based solutions excel at flight-mode or type identification [AlSaad2019, Allahham2020], but radio congestion and potential adversarial attacks are critical hurdles [Elleuch2024jamming, wang2024survey]. (2) Audio-based methods achieve high accuracy in controlled settings yet suffer from significant performance drops in high-noise or windy environments [Bernardini2017, Jeon2017, al2019droneaudio]. (3) Radar-based pipelines bypass visual occlusions but incur higher costs and calibration demands [mendis2016deep, wang2024survey]. (4) Vision-based approaches generally provide fine-grained recognition but degrade with lighting or weather changes [Wisniewski2022, Unlu2019]. (5) Multi-modal fusion can mitigate single-sensor weaknesses and reduce false positives [deng2024multi, Frid2024] yet raises complexity in real-time deployment [yuan2024mmaud]. Additionally, adversarial robustness remains an open problem, particularly where sensors or neural networks may be fooled by spoofing or deceptive signals [Medaiyese2021, Katta2022]. A continuing research direction involves efficient, low-latency architectures [xiao2025tame, wang2024survey] that can handle on-board or edge-based processing for rapid UAV detection and classification.
@@
| Jiang et al. (2023) [anti_uav] | RGB, IR | GlobalTrack + DFSC | Anti-UAV [anti_uav] | Released 318-sequence RGB-T UAV tracking benchmark. Achieved the best state accuracy score of 74.33% on the validation set and 65.41% on the test set |
| Wu et al. (2024) [Wu2024EagleEye] | RGB | DCF + dual-fovea + contrast + opponent-color (EECF) | Anti-UAV [anti_uav], USC-Drone [USC_Drone] | Real-time EECF tracker (86 FPS) via dual-fovea design. Achieved the state accuracy score of 66.42% on Anti-UAV, 70.45% on USC-Drone. |
| Wang et al. (2024) [wang2024target] | RGB | YOLOX + Swin-T + -IoU + BYTE | UAVSwarm [wang2022uavswarm] | Detection-driven approach with -IoU + ByteTrack. Achieved 78.2% tracking accuracy on UAVSwarm test split. |
| Deng et al.(2024) [deng2024multi] | RGB, RA, L | YOLOv9 + EfficientNet-B7 + LSTM | MMAUD [yuan2024mmaud] | Intergrated YOLOv9 + EfficientNet-B7 + LSTM/Kalman for classification and tracking. Achieved 2.21 Pose MSE Loss on 3D tracking of MMAUD and ranked 1st in UG2+ Challenge. |
@@
| Huang et al. (2021) [Huang2021SiamSTA] | IR | SiamSTA | 1st / 2nd Anti-UAV Challenges [anti_uav] [zhao20212nd] | Achieved 74.46 and 67.30 scores on 1st / 2nd Anti-UAV test-dev sets, ranking 1st. Extended SiamSTA with spatio-temporal attention + motion estimation. |
| Li et al. (2022) [li2022dual] | IR | SuperDiMP (ResNet50) | 2nd Anti-UAV Challenge [zhao20212nd] | Used SuperDiMP with dual-branch IR fusion. Achieved +5.84% SR and +4.52% PR over baseline. Focused on contrast inversion robustness. |
| Chen et al. (2022) [Chen2022SiamSTA] | IR | SiamSTA + CDCF | 1st Anti-UAV Challenge [anti_uav] | SiamSTA with CDCF and 3-stage redetection. Achieved top precision in 2nd Anti-UAV Challenge. Achieved 68.26 and 76.11 scores on Anti-UAV test and validation sets. Robust on fast/small UAVs. |
| Zhang et al. (2023) [zhang2023modality] | RGB, IR | SiamFusion | Anti-UAV [anti_uav] | Achieved overall 63.37 score on Anti-UAV with 14 FPS. Developed a dual-fusion tracker with modality + decision-level fusion. Used Swin backbone + Local-Global Converter. |
| Xie et al. (2023) [xie2023stftrack] | IR | FPN-based Siamese + STG-RPN + ID-RCNN (STFTrack) | Anti-UAV [anti_uav] | STFTrack using STG-RPN and ID-RCNN. Achieved 91.2% precision, 66.6% success @ 12.4 FPS. Addressed distractors and thermal crossover. |
| Zhang et al. (2024) [zhang2024precision] | IR | OSTrack + MCJT (RMCM + FSRM + ATUS ) | 1st / 2nd / 3rd Anti-UAV Challenges [anti_uav] [zhao20212nd] [zhao20233rd] | Achieved 69.65% Acc, 80.45% Acc, 65.88% Acc and 50.28% Acc on the 1st Anti-UAV test set, 1st Anti-UAV test-dev, 2nd Anti-UAV test-dev, and 3rd Anti-UAV validation set. Proposed MCJT (RMCM + FSRM + ATUS) with adaptive template and multi-consistency. Ran at 44 FPS. |
| Huang et al. (2024) [Huang2023SiamSRT] | IR | SiamSRT + C-C RPN + S-L RCNN + TMB + SCFD | Anti-UAV [anti_uav] | Achieved 71.64 and 80.04 scores on Anti-UAV testing and validation sets. Global Siamese tracker using Swin-T + consistency modules (TMB + SCFD). Avoided template degradation; outperformed multiple SOTA trackers. |
| Huang et al. (2024) [Huang2023AntiUAV410] | IR | SiamDT (Swin Transformer + Dual-Semantic RPN + Versatile R-CNN) | Anti-UAV410 [Huang2023AntiUAV410] | Introduced Anti-UAV410 dataset. Proposed SiamDT with dual semantic reasoning + background suppression. Achieved 68.19 and 71.65 scores on its testing and validation sets. |
@@
| Cheng et al. (2022) [Cheng2022AntiUAVTracking] | IR | Multi-Hybrid Attention + SiamRPN++ + YOLOv5 (SiamAD) | Anti-UAV [anti_uav] | SiamAD with CSAM/CAM + hierarchical discriminator + YOLOv5. 88.4% DP @ 37.1 FPS. Improved long-term IR tracking. |
| Yu et al. (2023) [Yu2023UnifiedTransformer] | IR | Transformer + YOLOv5 + LoFTR + GMM + OSTrack | 1st / 2nd Anti-UAV Challenges [anti_uav] [zhao20212nd] | Achieved 2nd place in Anti-UAV Challenge (3rd) with 68.8 score. And 75.13% and 80.38% accuracy on the 1st and 2nd test dev, individually. Integrated OSTrack, YOLOv5, LoFTR, GMM for multi-stage IR tracking. |
@@ 4 UAV Detection Approaches
(1) Radar-Based Detection. Radar methods detect UAVs by exploiting micro-Doppler signatures, providing resilience under low-visibility or harsh weather conditions [zhao2021mdrone, wang2024survey]. For instance, Zhao et al. [zhao2021mdrone] achieved decimeter-level localization at 10 Hz by combining CNNs with LSTMs, outperforming conventional radar baselines. However, radar sensors can suffer from interference and clutter, particularly in urban areas or heavily congested electromagnetic environments [Wu2024multi].
@@
(2) Vision-Based Detection. Vision-based techniques remain widely adopted, benefiting from readily available cameras and sophisticated CNN-based detectors [zhu2020vision, tan2020efficientdet]. Early work centered on region-based methods [ren2016faster] and YOLO architectures [bochkovskiy2020yolov4, wang2022yolov7], gradually evolving into more lightweight or attention-oriented variants [Cheng2024LocalGlobal, huang2024EDGSYOLOv8, hao2025yolov8]. While mosaic augmentation and specialized modules can improve small-object detection [Dadboud2021yolov5, Cheng2024LocalGlobal], vision-based approaches remain vulnerable to occlusions, clutter, and extreme distances. Researchers have thus explored thermal and infrared (IR) imaging [Sun2024multiyolov8, Nair2024ThermalFL] or speed-accuracy trade-offs for real-time embedded scenarios [bo2024yolov7gs, Elsayed2024LERFNetAE].
@@
(3) Multi-Modal Fusion. Single-modality approaches often fail under suboptimal conditions—low-visibility (vision), interference (radar), or noise (audio) [al2019droneaudio, uddin2022tvddt]. Consequently, multi-modal fusion has gained traction as a means of leveraging complementary sensor data [svanstrom2021dataset, arsenos2024]. Svanström et al. [svanstrom2021dataset] showed that augmenting YOLO-based RGB or thermal imaging with audio features increased F1 scores and reduced false alarms. Similarly, Wu et al. [Wu2024multi] fused radar and RGB streams via attention-based spatiotemporal networks to improve occlusion handling. Practical adoption, however, demands meticulous sensor synchronization, balanced hardware costs, robust computational efficiency, and real-time data association [Rebbapragada2024C2FDrone].
@@
Key Insights & Challenges. From prior work [wang2024survey, zhu2020vision, tan2020efficientdet], several recurring themes emerge. 1) Small UAV Detection: Limited pixel coverage and subtle micro-Doppler signatures complicate detection [He2024FogUAV, carion2020end, liu2021swin], prompting multi-scale feature extraction, deformable convolutions, and transformer backbones [tan2020efficientdet, Cheng2024LocalGlobal, huang2024EDGSYOLOv8]. Ensuring real-time operation at high accuracy remains challenging [wang2024survey, wang2022yolov7]. 2) Environmental Factors: Ambient noise (audio), lighting (vision), and radio interference (radar) degrade performance [al2019droneaudio, uddin2022tvddt, singh2024VisionUAV, zhao2021mdrone], highlighting the need for advanced clutter suppression and robust sensor hardware. 3) Generalization vs. Specialization: YOLO-based models offer generalizable frameworks [Dadboud2021yolov5, Cheng2024LocalGlobal, hao2025yolov8], but specialized domains (e.g., stealth UAVs) may require niche architectures or domain-adaptive strategies [Wu2024multi, bo2024yolov7gs]. 4) Multi-Modal Fusion: Complementary sensors effectively reduce false positives and enhance accuracy [svanstrom2021dataset, Wu2024multi], although issues like sensor alignment, real-time latency, and hardware cost must be resolved [arsenos2024, Rebbapragada2024C2FDrone]. 5) Benchmarking & Deployment: Heterogeneous data sources, labeling formats, and test conditions [du2018unmanned, Paweczyk2020RealWO, zhu2020vision] complicate standardized evaluations. Though laboratory demonstrations show promise, real-time performance in crowded outdoor settings remains largely underexplored [Fan2021DesignAI, Filkin2021UnmannedAV]. Overcoming these challenges will be crucial for field-ready anti-UAV systems.